# Efficient Quantification of Time-Series Prediction Error: Optimal Selection Conformal Prediction

### v1

@@ I Introduction
Uncertainty quantification using Bayesian methods include Bayesian Inference, Bayesian Neural Network and other variants [1, 2, 3]. Alternative approaches include concentrations bounds, such as Chernoff-Hoeffding (e.g., [4]), Clopper-Pearson (e.g., [5]). A more detailed review can be found in Section 2.3 of the survey paper [6]. However, Bayesian methods do not have finite-sample guarantees and become computationally intractable for large-scale problems; concentration bounds can be conservative in uncertainty quantification. An alternative to these methods is conformal prediction, which are presented next.
@@
Conformal Prediction (CP) constitutes a statistical framework that has been introduced [7] as a distribution-free and statistically rigorous tool to predict a $(1-\epsilon)$-confidence region for the trained prediction model, with $\epsilon\in(0,1)$. Under mild assumptions on a calibration dataset, CP provides tight finite-sample guarantees on both marginal coverage and conditional coverage. Although closely related with scenario optimization [8], a tool that has also been used for tight uncertainty quantification [9], CP focuses on a different aspect and thus complements the scenario approach (see [10] for some connections between CP and the scenario approach). Thanks to its simplicity, flexibility, and computational-efficiency, CP has been applied in probabilistic safe control synthesis and verification, such as moving-objects avoidance control [11, 12, 13, 14], probabilistic reachability analysis [15, 16, 17], probabilistic reachable sets construction [18, 19, 20].
However, while the confidence region produced by CP comes with rigorous probabilistic guarantees, its performance (e.g., size and shape of the region) still depends critically on one of its core components—the nonconformity measure (also called score function). Designing an appropriate score function for time-series uncertainty quantification is non-trivial. Existing works [21, 22, 23, 24] either suffer from 1) overly conservative confidence regions [21], 2) overfitting or fitting-errors [22, 23], or 3) computational intractability issues [23, 24]. To the best of our knowledge, no work seems to alleviate these issues at the same time.
In this paper, we aim to overcome these challenges when using CP for time series, thus opening the road for its use in multi-stage optimization and safety problems that require uncertainty quantification over entire trajectories rather than single time-steps. We propose a new parameterized score function that can be optimized to provide minimal-average-radius CP regions. Our CP method generates norm-ball regions, that are convex and as we will show also tight, for multi-dimensional time series and exhibits lower computational requirements compared to other algorithmic alternatives. Our proposed approach is directly applicable to control problems such as safe learning-based MPC [11, 12, 13, 14] and multi-stage safety verification [25].
@@
    We evaluate the efficacy of our approach numerically on 4 case studies and show that it produces valid conformal regions with the smallest size among baselines [21, 22, 23, 24, 26]. Specifically, the results suggest that our proposed approach reduces the conformal set size by 16.03%, 14.32%, 14.01%, 16.93% on the 4 case studies, respectively, compared to the previous State-of-the-Art (SOTA) method [24].
@@
    Our optimization program runtime requirements are mild; compared to previous SOTA [24], it leads to 8812.0, 78622.0, 14.4, 22.1 times faster computation on the 4 benchmark studies we have investigated numerically, respectively.
@@ II-A Problem Setting
Definition 2 (Validity, [7]).
@@ II-B Conformal Prediction and ICP Framework
Conformal Prediction (CP) is a model-agnostic and distribution-free tool that aims to quantify prediction uncertainty and produce valid prediction sets $\Gamma^{\epsilon}$ without assumptions on the data distribution. The most commonly used CP methods employ a computation-friendly framework called Inductive-Conformal-Prediction (ICP, see [7]), sometimes also called Split-Conformal-Prediction. Such framework assumes that the data in calibration set is exchangeable with the test data. Then we use a score function (non-conformity measure) $\mathcal{A}$ to assign each calibration data with a non-conformity score $R_{i}$, and find the $p$th smallest score with $p=\lceil(1-\epsilon)(N+1)\rceil$. Then any data point with score smaller or equal to this $p$th smallest score is in the valid conformal region.
@@
Definition 3 (Efficiency, [7]).
@@ III Optimal Selection Conformal Prediction (OSCP)
In this paper, we propose a new CP method for time series using the Re-calibrate ICP framework [22, 23, 24], which splits the $D_{\mathrm{cal}}$ into 2 halves, one for learning the parameters in a parameterized score function and another for calibration (as in the standard CP). We propose a novel parameterization that provides very tight conformal regions in theory and achieves SOTA on the experiments in Section IV. We also show that the computation of this method is much more efficient than the previous SOTA method.
@@ III-B OSCP: Algorithm Description, Validity and Efficiency
Proof: The construction of score function $\mathcal{A}$ depends only on $D_{\mathrm{cal},1}$, and does not include information from $D_{\mathrm{cal},2}$. Thus, the non-conformity score for each data in $D_{\mathrm{cal},2}$ is exchangeable with that of a new data drawn from $\mathcal{D}$. Let $p_{2}=\lceil(1-\epsilon)(n_{2}+1)\rceil$, by Lemma 1 in [27], the conformal regions defined as $\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1}^{(\mathrm{new})}):=\left\{\mathbf{Y}\in\mathbb{R}^{d\times T}\mid\mathcal{A}(\mathbf{Y}-\hat{\mathbf{Y}}^{(\mathrm{new})}_{0:T-1})\leq R^{[p_{2}]}\right\}$ has property that ${\mathbb{P}}\left(\mathbf{Y}^{(\mathrm{new})}_{0:T-1}\in\Gamma^{\epsilon}(\hat{\mathbf{Y}}^{(\mathrm{new})}_{0:T-1})\right)\geq 1-\epsilon$. Now, for any $\mathbf{Y}\in\mathbb{R}^{d\times T}$,
@@
The shape of CP regions that this method produces depends on the user pre-defined norm for calculating the absolute residuals. This can be viewed as a hyper-parameter for the method. For instance, $l_{2}$-norm produces ball-shaped regions, $l_{1}$ or $l_{\infty}$-norm induces hyper-rectangle regions, and positive-definite matrix $A$-norm results in ellipsoid-shaped regions (see [28]). Ellipsoidal regions is flexible, but may lead to over-fitting when the data is not enough to reflect its shape in high dimensional spaces. For the case that the dataset is small and we don’t have prior knowledge of the data distribution, we can assume the prediction residual follows a gaussian error and employ $l_{2}$-norm.
@@ IV Numerical Experiments
To demonstrate the performance of our method, we test on both simulated and real time-series with different time horizons $T$, dimensions $d$, and calibration dataset sizes $N$, taken from [22]. We compare our method with 5 baseline uncertainty-quantification (UQ) methods, and the results show that among all valid alternatives, the efficiency of our proposed approach outperforms the state-of-the-art method on all case studies.
@@
We selected MC-dropout [26] and CF-RNN [21] as the baseline approaches for Bayesian UQ method and CP for time series, respectively; as well as three recent approaches in CP for time series, namely, CopulaCPTS [22] and LCP [24] as parameter optimization methodologies, and CRD [23] as a convex CP baseline. Since both CRD and our method allow users to specify shapes, we test hyper-rectangle & ellipsoid shapes of CRD (denoted as CRD-Rect & CRD-Ell, respectively) and ball & ellipsoid shapes of our method ($l_{2} & OSCP$-Ell).
@@ IV-A Synthetic Datasets
According to [22], the first two datasets are generated from the Interacting Particle System [29], and extra Gaussian noises with standard deviations $\sigma=0.01$ and $0.05$ are added to the dynamics of particle simulations in two datasets, respectively. For each case, data is split into 2000/500/500 for training, calibration, testing, respectively. A prediction model is trained to predict the future dynamics $\mathbf{Y}\in\mathbb{R}^{d\times T}$ given the past observations $\mathbf{X}\in\mathbb{R}^{d\times\tau}$ of the particle simulation with $\tau=35,\ T=25,\ d=2$. Then all UQ methods are evaluated according to the procedures stated in Appendix IX. For UQ methods that require a further split of calibration dataset, the split ratio is set to be $0.5/0.5$ for $D_{\mathrm{cal},1}$ & $D_{\mathrm{cal},2}$ except LCP. We note that the optimization program in LCP becomes computationally intractable for this split, so we adopt $0.1/0.9$ split-ratio for LCP.
@@
The drone trajectory dataset is generated from [30] with added Gaussian noise of $\sigma=0.02$. The data-split is $600/200/200$ for training/calibration/testing. The prediction model forecasts a drone’s future trajectory given its past observations ($\tau=60,\ T=10,\ d=3$). After training the prediction model, all UQ methods are evaluated in a similar manner. For methods requiring a split of the calibration dataset, a split-ratio of $0.5/0.5$ is adopted (including LCP).
@@ IV-B Real Dataset: Covid-19 Daily Cases
We also conduct a case study on the real dataset, UK Covid-19 Daily Cases. We use the preprocessed dataset from [22]. [^2] The original Covid-19 dataset can be download at https://coronavirus.data.gov.uk/ Each time-series data in the preprocessed Covid-19 dataset corresponds to $150$-day daily cases from mid-September $2020$ to mid-February $2021$ at a region in UK. Then, $500/160/80$ time-series data are used for training/calibration/testing. The prediction model takes $100$ days of data as input, and outputs the subsequent $50$ days, i.e., $\tau=100,T=50,d=1$. For UQ methods that require a further split of the calibration data, the split ratio was set to $0.5/0.5$.
@@ IV-C Numerical Results
Our method outperforms all the baselines on 4 case studies for all 10 confidence levels ($0.5\%$ to $0.95\%$). Specifically, compared to previous SOTA method LCP [24] (the one with the smallest total volume among baselines that have empirical coverage no smaller than target ones), our method with $l_{2}$-norm, OSCP-$l_{2}$, reduces the total confidence region size (on average) of 16.03%, 14.32%, 14.01%, 16.93% on dataset Particle ($\sigma=0.01$), Particle ($\sigma=0.05$), Drone, Covid-19, respectively. When using an ellipsoidal confidence region, our method achieves further reductions in the region size (see Appendix IX-B).
@@
We also compare the runtime used in solving the optimization problem between our method and the previous SOTA (LCP, [24]). For the Particle ($\sigma=0.05$) dataset, LCP sometimes reach the time limit and terminate the simulation at $10000$s, so the actual computing time is higher than the reported result. In Table I, we can see that when target confidence is set as 90%, our method is 8812.0, 78622.0, 14.4, 22.1 times faster than LCP on the four datasets, respectively.
TABLE I: Comparison of optimization runtime (in sec) for target confidence $90\%$
Case Study Previous SOTA [24] OSCP-$l_{2}$
@@ IX-A Additional Information of Experiment Setup and Evaluation Metrics
All experiments are conducted on Windows 11 (64-bit) machine with i9-13900HX CPU (24 physical cores), 32GB RAM. The computations (including prediction models training) use CPU only. For solving the optimization problems in our method and in [24], we use the commercial solver Gurobi Optimizer (version 12.0.0 build v12.0.0rc1).
@@
Since we use the similar experiment settings (e.g., datasets, evaluation metrics, etc.) as in [22], we also use the experiment code of [22] to train the underlying time-series prediction model for all 4 case studies. [^3] Experiment code of [22] can be download at https://github.com/Rose-STL-Lab/CopulaCPTS Here is a brief description of the forecasting model $f_{\mathrm{pred}}$ for 4 case studies.
@@ IX-B Detailed Experiment Results and Visualizations
The complete results of our method with $l_{2}$-norm (OSCP-$l_{2}$) & with ellipsoidal norm (OSCP-Ell) and of all 5 baselines (LCP [24], CopulaCPTS [22], CRD [23], CF-RNN [21], MC-dropout [26]) on all 4 case studies are shown in Table II, Figure 3 & Figure 4.
In Table II, the data in bold have better efficiency (smaller total volume) than previous SOTA (LCP, [24]) and have coverages higher or equal to the target confidences, while the data in gray mean the empirical coverages are lower than the target confidences. It is important to note that if the empirical coverage of a method is slightly lower than the target confidence level (e.g. difference of coverages $\leq 2\%$), it does NOT necessarily mean that this method is invalid. The small gap in coverage can also be caused by numerical inaccuracy in computation (e.g., matrix inverse, eigen-value computations), or simply the randomness of empirical coverage (which converges to $\lceil(1-\epsilon)\cdot(N_{cal,2}+1)\rceil/(N_{cal,2}+1)$ in probability as $|D_{test}|\rightarrow\infty$, by the weak law of large numbers). However, to ensure a fair comparison of efficiency, here we only compare the results whose empirical coverages are no smaller than the target ones.
Figure 3: Case studies: Particle Datasets. The dashed reference line denotes the target confidences, and only methods with coverage curves at or above this line achieve the target coverages. The shaded region of each curve is the $\pm$ 1 standard error region.
Figure 4: Case studies: Drone Dataset & Covid-19 Dataset. The dashed reference line denotes the target confidences, and only methods with coverage curves at or above this line achieve the target coverages. The shaded region of each curve is the $\pm$ 1 standard error region.
@@
LCP [24] 85.4 $\pm$ 2.2 0.69 $\pm$ 0.07 90.1 $\pm$ 2.3 0.81 $\pm$ 0.08 95.5 $\pm$ 1.4 1.05 $\pm$ 0.12
CopulaCPTS [22] 85.3 $\pm$ 3.3 0.74 $\pm$ 0.11 89.8 $\pm$ 3.0 0.91 $\pm$ 0.3 94.4 $\pm$ 2.1 1.7 $\pm$ 1.16
CRD-Ell [23] 85.3 $\pm$ 3.0 0.93 $\pm$ 0.23 90.2 $\pm$ 2.4 1.02 $\pm$ 0.22 95.6 $\pm$ 1.4 1.24 $\pm$ 0.19
CRD-Rect [23] 92.7 $\pm$ 4.2 1.52 $\pm$ 0.41 94.5 $\pm$ 3.3 1.55 $\pm$ 0.41 96.1 $\pm$ 2.2 1.59 $\pm$ 0.39
CF-RNN [21] 97.0 $\pm$ 1.7 1.19 $\pm$ 0.22 98.3 $\pm$ 1.5 2.04 $\pm$ 1.03 99.4 $\pm$ 0.7 4.12 $\pm$ 1.46
@@
LCP [24] 85.4 $\pm$ 2.4 4.43 $\pm$ 0.29 90.5 $\pm$ 1.8 5.07 $\pm$ 0.3 95.5 $\pm$ 1.4 6.41 $\pm$ 0.67
CopulaCPTS [22] 85.6 $\pm$ 2.8 4.44 $\pm$ 0.46 90.0 $\pm$ 2.6 5.43 $\pm$ 0.86 93.4 $\pm$ 2.2 6.46 $\pm$ 1.28
CRD-Ell [23] 85.6 $\pm$ 2.8 4.88 $\pm$ 0.44 90.5 $\pm$ 2.0 5.5 $\pm$ 0.58 95.6 $\pm$ 1.1 7.11 $\pm$ 0.71
CRD-Rect [23] 92.0 $\pm$ 3.0 7.09 $\pm$ 0.82 93.2 $\pm$ 2.4 7.3 $\pm$ 1.05 95.4 $\pm$ 1.8 7.9 $\pm$ 1.06
CF-RNN [21] 92.6 $\pm$ 1.5 5.2 $\pm$ 0.29 95.0 $\pm$ 1.5 6.08 $\pm$ 0.67 97.6 $\pm$ 1.1 8.41 $\pm$ 1.29
MC-dropout [26] 33.8 $\pm$ 4.8 1.81 $\pm$ 0.07 42.7 $\pm$ 5.1 2.36 $\pm$ 0.09 55.5 $\pm$ 5.2 3.35 $\pm$ 0.13
@@
LCP [24] 84.4 $\pm$ 4.5 8.51 $\pm$ 2.09 89.9 $\pm$ 3.2 9.2 $\pm$ 2.29 95.2 $\pm$ 2.9 11.91 $\pm$ 4.83
CopulaCPTS [22] 87.8 $\pm$ 5.0 9.94 $\pm$ 1.92 91.5 $\pm$ 4.2 10.79 $\pm$ 2.07 95.6 $\pm$ 2.6 12.74 $\pm$ 1.95
CRD-Ell [23] 84.7 $\pm$ 3.7 163.64 $\pm$ 161.07 90.0 $\pm$ 3.1 207.23 $\pm$ 221.37 95.1 $\pm$ 2.1 305.0 $\pm$ 537.19

### v2

@@ I Introduction
Uncertainty quantification using Bayesian methods include Bayesian Inference, Bayesian Neural Network and other variants [fortunato2017bayesian, neal2011mcmc, kingma2013auto]. Alternative approaches include concentrations bounds, such as Chernoff-Hoeffding (e.g., [legay2019statistical]), Clopper-Pearson (e.g., [wang2019statistical]). A more detailed review can be found in Section 2.3 of the survey paper [lindemann2024formal]. However, Bayesian methods do not have finite-sample guarantees and become computationally intractable for large-scale problems; concentration bounds can be conservative in uncertainty quantification. An alternative to these methods is conformal prediction, which are presented next.
@@
Conformal Prediction (CP) constitutes a statistical framework that has been introduced [vovk2005algorithmic] as a distribution-free and statistically rigorous tool to predict a $(1-\epsilon)$-confidence region for the trained prediction model, with $\epsilon\in(0,1)$. Under mild assumptions on a calibration dataset, CP provides tight finite-sample guarantees on both marginal coverage and conditional coverage. Although closely related with scenario optimization [campi2018introduction], a tool that has also been used for tight uncertainty quantification [margellos2014road], CP focuses on a different aspect and thus complements the scenario approach (see [o2025bridging] for some connections between CP and the scenario approach). Thanks to its simplicity, flexibility, and computational-efficiency, CP has been applied in probabilistic safe control synthesis and verification, such as moving-objects avoidance control [lindemann2023safe-a, stamouli2024recursively, tonkens2023scalable, yu2023signal], probabilistic reachability analysis [bortolussi2019neural, cairoli2021neural, cairoli2023conformal], probabilistic reachable sets construction [hashemi2023data, hashemi2024statistical, tebjou2023data].
However, while the confidence region produced by CP comes with rigorous probabilistic guarantees, its performance (e.g., size and shape of the region) still depends critically on one of its core components—the nonconformity measure (also called score function). Designing an appropriate score function for time-series uncertainty quantification is non-trivial. Existing works [stankeviciute2021conformal, sun2024copula, Tumu2024multi-modal, cleaveland2024conformal] either suffer from 1) overly conservative confidence regions [stankeviciute2021conformal], 2) overfitting or fitting-errors [sun2024copula, Tumu2024multi-modal], or 3) computational intractability issues [Tumu2024multi-modal, cleaveland2024conformal]. To the best of our knowledge, no work seems to alleviate these issues at the same time.
In this paper, we aim to overcome these challenges when using CP for time series, thus opening the road for its use in multi-stage optimization and safety problems that require uncertainty quantification over entire trajectories rather than single time-steps. We propose a new parameterized score function that can be optimized to provide minimal-average-radius CP regions. Our CP method generates norm-ball regions, that are convex and as we will show also tight, for multi-dimensional time series and exhibits lower computational requirements compared to other algorithmic alternatives. Our proposed approach is directly applicable to control problems such as safe learning-based MPC [lindemann2023safe-a, stamouli2024recursively, tonkens2023scalable, yu2023signal] and multi-stage safety verification [Lindemann2023Conformal-b].
@@
    We evaluate the efficacy of our approach numerically on 4 case studies and show that it produces valid conformal regions with the smallest size among baselines [stankeviciute2021conformal, sun2024copula, Tumu2024multi-modal, cleaveland2024conformal, gal2016dropout]. Specifically, the results suggest that our proposed approach reduces the conformal set size by 16.03%, 14.32%, 14.01%, 16.93% on the 4 case studies, respectively, compared to the previous State-of-the-Art (SOTA) method [cleaveland2024conformal].
@@
    Our optimization program runtime requirements are mild; compared to previous SOTA [cleaveland2024conformal], it leads to 8812.0, 78622.0, 14.4, 22.1 times faster computation on the 4 benchmark studies we have investigated numerically, respectively.
@@ II-A Problem Setting
Definition 2 (Validity, [vovk2005algorithmic]).
@@ II-B Conformal Prediction and ICP Framework
Conformal Prediction (CP) is a model-agnostic and distribution-free tool that aims to quantify prediction uncertainty and produce valid prediction sets $\Gamma^{\epsilon}$ without assumptions on the data distribution. The most commonly used CP methods employ a computation-friendly framework called Inductive-Conformal-Prediction (ICP, see [vovk2005algorithmic]), sometimes also called Split-Conformal-Prediction. Such framework assumes that the data in calibration set is exchangeable with the test data. Then we use a score function (non-conformity measure) $\mathcal{A}$ to assign each calibration data with a non-conformity score $R_{i}$, and find the $p$th smallest score with $p=\lceil(1-\epsilon)(N+1)\rceil$. Then any data point with score smaller or equal to this $p$th smallest score is in the valid conformal region.
@@
Definition 3 (Efficiency, [vovk2005algorithmic]).
@@ III Optimal Selection Conformal Prediction (OSCP)
In this paper, we propose a new CP method for time series using the Re-calibrate ICP framework [sun2024copula, Tumu2024multi-modal, cleaveland2024conformal], which splits the $D_{\mathrm{cal}}$ into 2 halves, one for learning the parameters in a parameterized score function and another for calibration (as in the standard CP). We propose a novel parameterization that provides very tight conformal regions in theory and achieves SOTA on the experiments in Section IV. We also show that the computation of this method is much more efficient than the previous SOTA method.
@@ III-B OSCP: Algorithm Description, Validity and Efficiency
Proof: The construction of score function $\mathcal{A}$ depends only on $D_{\mathrm{cal},1}$, and does not include information from $D_{\mathrm{cal},2}$. Thus, the non-conformity score for each data in $D_{\mathrm{cal},2}$ is exchangeable with that of a new data drawn from $\mathcal{D}$. Let $p_{2}=\lceil(1-\epsilon)(n_{2}+1)\rceil$, by Lemma 1 in [Tibshirani2019Conformal], the conformal regions defined as $\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1}^{(\mathrm{new})}):=\left\{\mathbf{Y}\in\mathbb{R}^{d\times T}\mid\mathcal{A}(\mathbf{Y}-\hat{\mathbf{Y}}^{(\mathrm{new})}_{0:T-1})\leq R^{[p_{2}]}\right\}$ has property that ${\mathbb{P}}\left(\mathbf{Y}^{(\mathrm{new})}_{0:T-1}\in\Gamma^{\epsilon}(\hat{\mathbf{Y}}^{(\mathrm{new})}_{0:T-1})\right)\geq 1-\epsilon$. Now, for any $\mathbf{Y}\in\mathbb{R}^{d\times T}$,
@@
The shape of CP regions that this method produces depends on the user pre-defined norm for calculating the absolute residuals. This can be viewed as a hyper-parameter for the method. For instance, $l_{2}$-norm produces ball-shaped regions, $l_{1}$ or $l_{\infty}$-norm induces hyper-rectangle regions, and positive-definite matrix $A$-norm results in ellipsoid-shaped regions (see [Xu2024Conformal]). Ellipsoidal regions is flexible, but may lead to over-fitting when the data is not enough to reflect its shape in high dimensional spaces. For the case that the dataset is small and we don’t have prior knowledge of the data distribution, we can assume the prediction residual follows a gaussian error and employ $l_{2}$-norm.
@@ IV Numerical Experiments
To demonstrate the performance of our method, we test on both simulated and real time-series with different time horizons $T$, dimensions $d$, and calibration dataset sizes $N$, taken from [sun2024copula]. We compare our method with 5 baseline uncertainty-quantification (UQ) methods, and the results show that among all valid alternatives, the efficiency of our proposed approach outperforms the state-of-the-art method on all case studies.
@@
We selected MC-dropout [gal2016dropout] and CF-RNN [stankeviciute2021conformal] as the baseline approaches for Bayesian UQ method and CP for time series, respectively; as well as three recent approaches in CP for time series, namely, CopulaCPTS [sun2024copula] and LCP [cleaveland2024conformal] as parameter optimization methodologies, and CRD [Tumu2024multi-modal] as a convex CP baseline. Since both CRD and our method allow users to specify shapes, we test hyper-rectangle & ellipsoid shapes of CRD (denoted as CRD-Rect & CRD-Ell, respectively) and ball & ellipsoid shapes of our method ($l_{2} & OSCP$-Ell).
@@ IV-A Synthetic Datasets
According to [sun2024copula], the first two datasets are generated from the Interacting Particle System [Kipf2018Neural], and extra Gaussian noises with standard deviations $\sigma=0.01$ and $0.05$ are added to the dynamics of particle simulations in two datasets, respectively. For each case, data is split into 2000/500/500 for training, calibration, testing, respectively. A prediction model is trained to predict the future dynamics $\mathbf{Y}\in\mathbb{R}^{d\times T}$ given the past observations $\mathbf{X}\in\mathbb{R}^{d\times\tau}$ of the particle simulation with $\tau=35,\ T=25,\ d=2$. Then all UQ methods are evaluated according to the procedures stated in Appendix IX. For UQ methods that require a further split of calibration dataset, the split ratio is set to be $0.5/0.5$ for $D_{\mathrm{cal},1}$ & $D_{\mathrm{cal},2}$ except LCP. We note that the optimization program in LCP becomes computationally intractable for this split, so we adopt $0.1/0.9$ split-ratio for LCP.
@@
The drone trajectory dataset is generated from [sakai2018pythonroboticspythoncodecollection] with added Gaussian noise of $\sigma=0.02$. The data-split is $600/200/200$ for training/calibration/testing. The prediction model forecasts a drone’s future trajectory given its past observations ($\tau=60,\ T=10,\ d=3$). After training the prediction model, all UQ methods are evaluated in a similar manner. For methods requiring a split of the calibration dataset, a split-ratio of $0.5/0.5$ is adopted (including LCP).
@@ IV-B Real Dataset: Covid-19 Daily Cases
We also conduct a case study on the real dataset, UK Covid-19 Daily Cases. We use the preprocessed dataset from [sun2024copula]. [^2] The original Covid-19 dataset can be download at https://coronavirus.data.gov.uk/ Each time-series data in the preprocessed Covid-19 dataset corresponds to $150$-day daily cases from mid-September $2020$ to mid-February $2021$ at a region in UK. Then, $500/160/80$ time-series data are used for training/calibration/testing. The prediction model takes $100$ days of data as input, and outputs the subsequent $50$ days, i.e., $\tau=100,T=50,d=1$. For UQ methods that require a further split of the calibration data, the split ratio was set to $0.5/0.5$.
@@ IV-C Numerical Results
Our method outperforms all the baselines on 4 case studies for all 10 confidence levels ($0.5$ to $0.95$). Specifically, compared to previous SOTA method LCP [cleaveland2024conformal] (the one with the smallest total volume among baselines that have empirical coverage no smaller than target ones), our method with $l_{2}$-norm, OSCP-$l_{2}$, reduces the total confidence region size (on average) of 16.03%, 14.32%, 14.01%, 16.93% on dataset Particle ($\sigma=0.01$), Particle ($\sigma=0.05$), Drone, Covid-19, respectively. When using an ellipsoidal confidence region, our method achieves further reductions in the region size (see Appendix IX-B).
@@
We also compare the runtime used in solving the optimization problem between our method and the previous SOTA (LCP, [cleaveland2024conformal]). For the Particle ($\sigma=0.05$) dataset, LCP sometimes reach the time limit and terminate the simulation at $10000$s, so the actual computing time is higher than the reported result. In Table I, we can see that when target confidence is set as $0.9$, our method is 8812.0, 78622.0, 14.4, 22.1 times faster than LCP on the four datasets, respectively.
TABLE I: Comparison of optimization runtime (in sec) for target confidence $1-\epsilon=0.9$
Case Study Previous SOTA [cleaveland2024conformal] OSCP-$l_{2}$
@@ IX-A Additional Information of Experiment Setup and Evaluation Metrics
All experiments are conducted on Windows 11 (64-bit) machine with i9-13900HX CPU (24 physical cores), 32GB RAM. The computations (including prediction models training) use CPU only. For solving the optimization problems in our method and in [cleaveland2024conformal], we use the commercial solver Gurobi Optimizer (version 12.0.0 build v12.0.0rc1).
@@
Since we use the similar experiment settings (e.g., datasets, evaluation metrics, etc.) as in [sun2024copula], we also use the experiment code of [sun2024copula] to train the underlying time-series prediction model for all 4 case studies. [^3] Experiment code of [sun2024copula] can be download at https://github.com/Rose-STL-Lab/CopulaCPTS Here is a brief description of the forecasting model $f_{\mathrm{pred}}$ for 4 case studies.
@@ IX-B Detailed Experiment Results and Visualizations
The complete results of our method with $l_{2}$-norm (OSCP-$l_{2}$) & with ellipsoidal norm (OSCP-Ell) and of all 5 baselines (LCP [cleaveland2024conformal], CopulaCPTS [sun2024copula], CRD [Tumu2024multi-modal], CF-RNN [stankeviciute2021conformal], MC-dropout [gal2016dropout]) on all 4 case studies are shown in Table II, Figure 3 & Figure 4.
In Table II, the data in bold have better efficiency (smaller total volume) than previous SOTA (LCP, [cleaveland2024conformal]) and have coverages higher or equal to the target confidences, while the data in gray mean the empirical coverages are lower than the target confidences. It is important to note that if the empirical coverage of a method is slightly lower than the target confidence level (e.g. difference of coverages $\leq 2\%$), it does NOT necessarily mean that this method is invalid. The small gap in coverage can also be caused by numerical inaccuracy in computation (e.g., matrix inverse, eigen-value computations), or simply the randomness of empirical coverage (which converges to $\lceil(1-\epsilon)\cdot(N_{cal,2}+1)\rceil/(N_{cal,2}+1)$ in probability as $|D_{test}|\rightarrow\infty$, by the weak law of large numbers). However, to ensure a fair comparison of efficiency, here we only compare the results whose empirical coverages are no smaller than the target ones.
Figure 3: Case studies: Particle Datasets. The dashed reference line denotes the target confidences, and only methods with coverage curves at or above this line achieve the target coverages. The shaded region of each curve is the $\pm$ 1 standard error region. In Area graphs, lower curves indicate better performance.
Figure 4: Case studies: Drone Dataset & Covid-19 Dataset. The dashed reference line denotes the target confidences, and only methods with coverage curves at or above this line achieve the target coverages. The shaded region of each curve is the $\pm$ 1 standard error region. In Length/Volume graphs, lower curves indicate better performance.
@@
LCP [cleaveland2024conformal] 85.4 $\pm$ 2.2 0.69 $\pm$ 0.07 90.1 $\pm$ 2.3 0.81 $\pm$ 0.08 95.5 $\pm$ 1.4 1.05 $\pm$ 0.12
CopulaCPTS [sun2024copula] 85.3 $\pm$ 3.3 0.74 $\pm$ 0.11 89.8 $\pm$ 3.0 0.91 $\pm$ 0.3 94.4 $\pm$ 2.1 1.7 $\pm$ 1.16
CRD-Ell [Tumu2024multi-modal] 85.3 $\pm$ 3.0 0.93 $\pm$ 0.23 90.2 $\pm$ 2.4 1.02 $\pm$ 0.22 95.6 $\pm$ 1.4 1.24 $\pm$ 0.19
CRD-Rect [Tumu2024multi-modal] 92.7 $\pm$ 4.2 1.52 $\pm$ 0.41 94.5 $\pm$ 3.3 1.55 $\pm$ 0.41 96.1 $\pm$ 2.2 1.59 $\pm$ 0.39
CF-RNN [stankeviciute2021conformal] 97.0 $\pm$ 1.7 1.19 $\pm$ 0.22 98.3 $\pm$ 1.5 2.04 $\pm$ 1.03 99.4 $\pm$ 0.7 4.12 $\pm$ 1.46
@@
LCP [cleaveland2024conformal] 85.4 $\pm$ 2.4 4.43 $\pm$ 0.29 90.5 $\pm$ 1.8 5.07 $\pm$ 0.3 95.5 $\pm$ 1.4 6.41 $\pm$ 0.67
CopulaCPTS [sun2024copula] 85.6 $\pm$ 2.8 4.44 $\pm$ 0.46 90.0 $\pm$ 2.6 5.43 $\pm$ 0.86 93.4 $\pm$ 2.2 6.46 $\pm$ 1.28
CRD-Ell [Tumu2024multi-modal] 85.6 $\pm$ 2.8 4.88 $\pm$ 0.44 90.5 $\pm$ 2.0 5.5 $\pm$ 0.58 95.6 $\pm$ 1.1 7.11 $\pm$ 0.71
CRD-Rect [Tumu2024multi-modal] 92.0 $\pm$ 3.0 7.09 $\pm$ 0.82 93.2 $\pm$ 2.4 7.3 $\pm$ 1.05 95.4 $\pm$ 1.8 7.9 $\pm$ 1.06
CF-RNN [stankeviciute2021conformal] 92.6 $\pm$ 1.5 5.2 $\pm$ 0.29 95.0 $\pm$ 1.5 6.08 $\pm$ 0.67 97.6 $\pm$ 1.1 8.41 $\pm$ 1.29
MC-dropout [gal2016dropout] 33.8 $\pm$ 4.8 1.81 $\pm$ 0.07 42.7 $\pm$ 5.1 2.36 $\pm$ 0.09 55.5 $\pm$ 5.2 3.35 $\pm$ 0.13
@@
LCP [cleaveland2024conformal] 84.4 $\pm$ 4.5 8.51 $\pm$ 2.09 89.9 $\pm$ 3.2 9.2 $\pm$ 2.29 95.2 $\pm$ 2.9 11.91 $\pm$ 4.83
CopulaCPTS [sun2024copula] 87.8 $\pm$ 5.0 9.94 $\pm$ 1.92 91.5 $\pm$ 4.2 10.79 $\pm$ 2.07 95.6 $\pm$ 2.6 12.74 $\pm$ 1.95
CRD-Ell [Tumu2024multi-modal] 84.7 $\pm$ 3.7 163.64 $\pm$ 161.07 90.0 $\pm$ 3.1 207.23 $\pm$ 221.37 95.1 $\pm$ 2.1 305.0 $\pm$ 537.19